OpenStudy (anonymous):

The slope of a line is -1/5, and the y-intercept is 5. What is the equation of the line written in general form? x - 5y - 25 = 0 x - 5y - 5 = 0 x + 5y - 25 = 0

Mehek (mehek14):

which one do you think?

Mehek (mehek14):

If you solve for y, the equation should come out like this \(y=-\dfrac{1}{5}x+5\)

OpenStudy (anonymous):

Okay, let me try that...give me a minute! ^_^

Mehek (mehek14):

alright

OpenStudy (anonymous):

Okay so, for fractions I multiply by common denominator correct?

Mehek (mehek14):

no you don't need a common denominator to multiply or divide fractions that's only for addition or subtraction

Mehek (mehek14):

Do you want me to hep you solve them?

OpenStudy (anonymous):

Which would make: y=-1/5x+5 y(5) = -1/5(5)x+5(5) 5 = -x+25 Yes?

OpenStudy (anonymous):

I just need to be walked through it...I get it once I understand fully! I don't want to be given answers.

OpenStudy (anonymous):

x+5y=25 x+5y-25=0 !!! Yes? :D